OpenStudy (anonymous):

write an equation in slope interecpt form for the line that passes through 4,7 and has a slope of -3/4

OpenStudy (bahrom7893):

y-7=(-3/4)(x-4) y-7=(-3/4)x+3 y=(-3/4)x+10

OpenStudy (anonymous):

much thanks!